A hydrogen x has relative molecular mass 56 consist of 87.8 percent by mass of carbon and 12.2 percent hydrogen determine the empirical formula of x and hence give its molecular formula and the iupac nomenclature of the compound formed.
1
Expert's answer
2019-03-26T04:22:36-0400
"w(C) = \frac {Ar(C)\times x}{Mr(C_xH_y)}"

"\therefore x = \frac {w(C) \times Mr(C_xH_y)}{Ar(C)} = \frac{0.878 \times 56}{12} = 4"

"w(H) = \frac {Ar(H)\times y}{Mr(C_xH_y)}"

"\therefore y = \frac {w(H) \times Mr(C_xH_y)}{Ar(H)} = \frac{0.122 \times 56}{1} = 6.832"

y should be integer. There obviously should be a mistake in the task as there is no such compound C4H7.

Mollar mass 56 g/mol corresponds to alkene C4H8, where w(C) = 85.7 % and w(H) = 14.3 %.

If so, we can determine x and y:


"x = \frac {w(C) \times Mr(C_xH_y)}{Ar(C)} = \frac{0.857 \times 56}{12} = 4"

"y = \frac {w(H) \times Mr(C_xH_y)}{Ar(H)} = \frac{0.143 \times 56}{1} = 8"

The molecular formula is "C_4H_8" butene-1 or butene-2 , empirical formula is "CH_2"